OpenStudy (anonymous):

Water must be mixed with a 15% hydrochloric acid solution to dilute the concentration to 5%. How many liters of water should be added if we need 19 liters of the 5% solution? Round to the nearest tenth. Answer is 12.7 liters..cannot figure out equation to get 12.7 liters...keep coming up with 6.33333..ect. Cannot figure out how to check problem, with answer.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

h+w = 19 ; h = 19-w h(.15) +w(1) = 19(.05) .......................................... (19-w)(.15) +w(1) = 19(.05) 2.85 -.15w +w = 0.95 .85w = -1.9 w = -2.235..now where did i go wrong lol

OpenStudy (amistre64):

is water 0 instead of 1?

OpenStudy (dumbcow):

yes water is 0% concentration

OpenStudy (amistre64):

(19-w)(.15) +w(0) = 19(.05) 2.85 -.15w = 0.95 -.15w = -1.9 w = 12.666......

OpenStudy (amistre64):

we got hard water here kol

OpenStudy (anonymous):

Your answer is very close, amistre64...

OpenStudy (anonymous):

if we rounded it, it would get me 12.7liters..

OpenStudy (dumbcow):

they prob rounded the answer

OpenStudy (amistre64):

w*1 lol .....

OpenStudy (anonymous):

Could either of you also explain how to check this problem with the answer?

OpenStudy (amistre64):

well, by substituting it back into the original equation to see if it works out

OpenStudy (amistre64):

amount(%) + amount(%) = total amount(%)

OpenStudy (dumbcow):

(12.7)(.05) = (6.3)(.15) +0

OpenStudy (dumbcow):

oops (19)(.05) = (6.3)(.15)

OpenStudy (anonymous):

I'll try amestre's equation now...if I don't come up with amestre's answer...I'll be coming back to this site definitely..maybe within the hour. :)

OpenStudy (anonymous):

amistre64, could you explain why you set up your equation the way you did, in relation to the problem? How does (19-w)(.15)+w(0) relate to the problem? Sorry for asking such dumb questions...but it's late and I relate to words better than just numbers..

OpenStudy (amistre64):

well.... we know that we want a mix of hydroacid and water that equals 19 liters; that restricts us to: h+w = 19. It is easier to solve an equation when everything is in one variable; so since we want the amount of water, lets get h in terms of w h = 19 liters - water amount (h=19-w)

OpenStudy (amistre64):

h(15%) is the strength of the hydroacid (19-w)(.15) is stateing the same thing but using terms in water...

OpenStudy (amistre64):

the rest should be self explanatory id assume

OpenStudy (anonymous):

I'll look at the equation again,see if I can get my lightbulb on...

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

At any rate, I've saved your formula in a document amistre64, in case I can't figure it out tonight...:)

OpenStudy (amistre64):

itll make more sense after a 4course meal :)

OpenStudy (anonymous):

I don't think I've ever had one of those...:P

OpenStudy (anonymous):

The light-bulb is starting to spark...but it's not on quite yet..

OpenStudy (anonymous):

Okay, listening to some of Bach's music has helped turn the light-bulb on...and now I'm wondering how it can be this simple...:P

OpenStudy (anonymous):

Thanks for your help, guys...I'll definitely be using this site more often from now on, with me majoring in nursing at the moment. :)

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

Hold on, where did you get the 2.85 in your equation, amistre64?

OpenStudy (anonymous):

dumbcow, do you know where amistre64's 2.85 came from, in his equation?

OpenStudy (anonymous):

Could somebody explain this equation step by step, all the way through to the end? Don't assume something is self-explanatory...it's never like that with me.